OpenStudy (anonymous):

f(x) = 2x3 + 3x2 − 432x Part (a) Find the intervals on which f is increasing. Find the interval on which f is decreasing. Part (b) Find the local minimum and maximum values of f. Part (c) Find the inflection point. Find the interval on which f is concave up. Find the interval on which f is concave down.

OpenStudy (bahrom7893):

a) find f': 6x^2 + 6x - 432 = 0 then find where its positive, that is when f is increasing.

OpenStudy (bahrom7893):

when its negative f will be decreasing, at 0 u will have local max and mins

OpenStudy (bahrom7893):

c) f" = 12x + 6 = 0 12x+6 = 0 and the signs of f" change before and after that 0, then u have an inflection point

OpenStudy (bahrom7893):

concave up - f" >0 concave down - f"<0, sorry i gtg just do all the steps above..

OpenStudy (anonymous):

Thanks a lot really appreciate it.